OpenStudy (anonymous):

what are the zeroes of the polynomial function f(x)=(x+4)(2x-1)(x-7)

OpenStudy (asnaseer):

what this is asking for is: for what values of x is f(x) equal to zero. now, if: \(a\times b\times c=0\)then either a=0 or b=0 or c=0 agreed?

OpenStudy (anonymous):

a=4,1/2,-7 b=-4,-1/2,7 c=4,-1/2,-7 d=-4,1/2/7 those are the options

OpenStudy (asnaseer):

I didn't ask you for the options

OpenStudy (anonymous):

well those are the options

OpenStudy (asnaseer):

if: \(a\times b\times c=0\) then either a=0 or b=0 or c=0 do you agree with this?

OpenStudy (anonymous):

yes i agree but if u had to choose one of the options which would it be because thats the way the question was asked

OpenStudy (asnaseer):

ok, so now if we set f(x)=0 we get:\[(x+4)(2x-1)(x-7)=0\]so here we have three items, namely (x+4), (2x-1) and (x-7) that are being multiplied together to give zero. so wat can you now deduce from this?